# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v12

@@ 5.1 Theorem of Model Interpretability
According to the view that interpretability should be a concept dependent on the domain of discourse and the corresponding logical system, and that interpretability should also be related to the specific inputs and outputs of the model [16], the following concepts are established based on the MLT-MF framework:
@@ 5.2 Theorem of Ethical Safety Assurance
By the maximum independent set theorem for finite hypergraphs[36], $H_{or}$ necessarily contains an independent set $G_{\max}$ of maximal cardinality. Suppose there exists another ethically safe subset $K$ with
@@ 5.3 Theorem on TVD Upper Bounds
Using KL-divergence to measure the difference between the true distribution and the hypothesized model distribution, we have [29]:
@@
Since both $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can use KL-divergence to estimate an upper bound on their TVD. By Pinsker’s inequality [30], we have:
@@
Distinct from traditional generalization theories such as VC dimension and Rademacher complexity, Theorem 3, based on the MLT-MF framework, utilizes $P_{Mq}$ and $|S_{oq}|-|S_{ou}\cap S_{oq}|$—which reflect the overlap between test data and the model’s original information—as well as the entropy $H(p_{q})$ of test data in the non-overlapping region, as evaluation metrics to provide an explicit quantitative estimation of the TVD. Since TVD reflects the degree of distributional alignment of the model, independent of specific models and tasks, it is generally applicable for analyzing training performance. Moreover, for classification tasks with 0-1 loss, the upper bound of the generalization error is equivalent to the TVD [31], while for regression tasks, the upper bound of the generalization error is given by the product of the TVD and the Lipschitz loss [32]. Therefore, TVD serves as an indicator for estimating a model’s generalization error. In this way, Theorem 3 offers a novel perspective and practical guidance for model design and training data acquisition.
@@ Appendix A Formal Representation of States
For simplicity, $\mathcal{L}$ does not define connectives such as “and” $\land$, “or” $\vee$, “if and only if” $\leftrightarrow$, or the quantifier “exists” $\exists$, and other commonly used logical symbols. However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of existing symbols and can make many formulas more concise [37]. Therefore, they are introduced as defined symbols in subsequent discussions.
@@ Appendix B Finite Automaton States
Using the formal expression of states, we can study in depth the special and important object of finite automata [38].
@@
Proof: Let $M=\left\langle Q,R,U,\delta,\lambda\right\rangle$ be a finite automaton, and $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$ the (strictly increasing) time points at which $M$ performs input, output, or state‐transition actions. By finite‐automaton theory [38], $Q=\left\{q_{t_{i}}\middle|i=1,\ldots,n-1\right\}$, $R=\left\{r_{t_{i}}\middle|i=2,\ldots,n\right\}$, and $U=\left\{u_{t_{i}}\middle|i=1,\ldots,n\right\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;

### v13

@@ 5.1 Theorem of Model Interpretability
According to the view that interpretability should be a concept dependent on the domain of discourse and the corresponding logical system, and that interpretability should also be related to the specific inputs and outputs of the model [glanois2024survey], the following concepts are established based on the MLT-MF framework:
@@ 5.2 Theorem of Ethical Safety Assurance
By the maximum independent set theorem for finite hypergraphs[Berge1973Graphs], $H_{or}$ necessarily contains an independent set $G_{\max}$ of maximal cardinality. Suppose there exists another ethically safe subset $K$ with
@@ 5.3 Theorem on TVD Upper Bounds
Using KL-divergence to measure the difference between the true distribution and the hypothesized model distribution, we have [goodfellow2016deep]:
@@
Since both $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can use KL-divergence to estimate an upper bound on their TVD. By Pinsker’s inequality [pinsker1964information], we have:
@@
Distinct from traditional generalization theories such as VC dimension and Rademacher complexity, Theorem 3, based on the MLT-MF framework, utilizes $P_{Mq}$ and $|S_{oq}|-|S_{ou}\cap S_{oq}|$—which reflect the overlap between test data and the model’s original information—as well as the entropy $H(p_{q})$ of test data in the non-overlapping region, as evaluation metrics to provide an explicit quantitative estimation of the TVD. Since TVD reflects the degree of distributional alignment of the model, independent of specific models and tasks, it is generally applicable for analyzing training performance. Moreover, for classification tasks with 0-1 loss, the upper bound of the generalization error is equivalent to the TVD [rodriguez2021tighter], while for regression tasks, the upper bound of the generalization error is given by the product of the TVD and the Lipschitz loss [devroye2013probabilistic]. Therefore, TVD serves as an indicator for estimating a model’s generalization error. In this way, Theorem 3 offers a novel perspective and practical guidance for model design and training data acquisition.
@@ Appendix A Formal Representation of States
For simplicity, $\mathcal{L}$ does not define connectives such as “and” $\land$, “or” $\vee$, “if and only if” $\leftrightarrow$, or the quantifier “exists” $\exists$, and other commonly used logical symbols. However, mathematical logic theory demonstrates that these symbols are entirely equivalent to appropriate combinations of existing symbols and can make many formulas more concise [hamilton1988logic]. Therefore, they are introduced as defined symbols in subsequent discussions.
@@ Appendix B Finite Automaton States
Using the formal expression of states, we can study in depth the special and important object of finite automata [editorial2002dictionary].
@@
Proof: Let $M=\left\langle Q,R,U,\delta,\lambda\right\rangle$ be a finite automaton, and $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$ the (strictly increasing) time points at which $M$ performs input, output, or state‐transition actions. By finite‐automaton theory [editorial2002dictionary], $Q=\left\{q_{t_{i}}\middle|i=1,\ldots,n-1\right\}$, $R=\left\{r_{t_{i}}\middle|i=2,\ldots,n\right\}$, and $U=\left\{u_{t_{i}}\middle|i=1,\ldots,n\right\}$ are nonempty finite sets of input symbols, output symbols, and states, respectively;